OpenStudy (anonymous):

anyone? =( A particle moves along the x axis. Its position is given by the equation x = 1.9 + 2.5t − 3.7t2 with x in meters and t in seconds. (a) Determine its position when it changes direction. ____m (b) Determine its velocity when it returns to the position it had at t = 0? (Indicate the direction of the velocity with the sign of your answer.) ____m/s

OpenStudy (anonymous):

changes direction when the derivative goes from being positive to negative or vice versa

OpenStudy (anonymous):

\[x(t)=1.9+2.5t-3.7t^2\] \[x'(t)=2.5-7.4t\] put \[2.5-7.4t=0\] get \[t=\frac{2.5}{7.4}\]

OpenStudy (anonymous):

we could have also done this by simply noting that this as a parabola open down so vertex is at \[-\frac{b}{2a}\] would have the same answer

OpenStudy (anonymous):

but if figured this was cacl class by the next question.

OpenStudy (anonymous):

at \[t=0\] you were at \[x(0)=1.9\] so set \[x(t)=1.9\] and solve for t

OpenStudy (anonymous):

then substitute that number back in to the derivative to get the velocity.

OpenStudy (anonymous):

stupid decimals. i get \[t=.676\] rounded and \[x'(.676)=2.4-7.4\times .676\] whatever that is

OpenStudy (anonymous):

gonna try it now =)